# Project Page: https://yikai-wang.github.io/seele ReS Dataset: https://github.com/Yikai-Wang/ReS Repositioning the Subject within Image

### v2

@@ Abstract
Current image manipulation primarily centers on static manipulation, such as replacing specific regions within an image or altering its overall style. In this paper, we introduce an innovative dynamic manipulation task, subject repositioning. This task involves relocating a user-specified subject to a desired position while preserving the image’s fidelity. Our research reveals that the fundamental sub-tasks of subject repositioning, which include filling the void left by the repositioned subject, reconstructing obscured portions of the subject and blending the subject to be consistent with surrounding areas, can be effectively reformulated as a unified, prompt-guided inpainting task. Consequently, we can employ a single diffusion generative model to address these sub-tasks using various task prompts learned through our proposed task inversion technique. Additionally, we integrate pre-processing and post-processing techniques to further enhance the quality of subject repositioning. These elements together form our SEgment-gEnerate-and-bLEnd (SEELE) framework. To assess SEELE’s effectiveness in subject repositioning, we assemble a real-world subject repositioning dataset called ReS. Results of SEELE on ReS demonstrate its efficacy.
Keywords:
Subject Repositioning Inpainting Completion
@@ 1 Introduction
In 2023, Google Photos introduced an AI editing feature allowing users to reposition subjects within their images [1]. However, a lack of technical documentation limits understanding of this feature. Some researches have touched on aspects of it. Iizuka et al. [23] explored object repositioning before the deep learning era, using user inputs like ground regions and bounding boxes. In the deep learning era, fields like scene decomposition [81] and de-occlusion [76] enable manipulation of object positions. This paper addresses general Subject Repositioning (SubRep) task without explicit scene understanding. Our aim is to address SubRep via a meticulously crafted solution, driven by a single diffusion model.
@@
From an academic perspective, this task falls under image manipulation [18, 24, 83, 66, 14, 15, 79]. Recent advancements in large-scale generative models have fueled interest in this field. These models, including generative adversarial models [20], variational autoencoders [32], auto-regressive models [65], and notably, diffusion models [60], demonstrate impressive image manipulation capabilities with expanding model architectures and training datasets [56, 28, 6]. However, current image manipulation methods primarily target "static" alterations, modifying specific image regions using cues like natural language, sketches, or layouts [14, 79, 15]. Another aspect involves style-transfer tasks, transforming overall image styles such as converting photos into anime pictures or paintings [7, 66, 25]. Some techniques extend to video manipulation, dynamically altering style or subjects over time [30, 70, 16]. In contrast, subject repositioning dynamically relocates selected subjects within a single image while leaving the rest unchanged.
@@
The SubRep task involves multiple stages, including non-generative and generative tasks. Existing pre-trained models are effective for non-generative tasks like segmenting subjects [33] and estimating occlusion relationships [54]. Our focus lies on the generative tasks of SubRep, including: i) Subject removal: The generative model must fill voids left after repositioning without introducing new elements. ii) Subject completion: If the repositioned subject is partially obscured, the model must complete it to maintain integrity. iii) Subject harmonization: The repositioned subject should blend with surrounding areas. All these sub-tasks demand unique generative capabilities.
@@
The most powerful text-to-image diffusion models [49, 21, 58, 53, 56] show potential promise for SubRep. However, a key challenge is finding suitable text prompts, as these models are usually trained with image captions rather than task-specific instructions. The best prompts are often image-dependent and hard to generalize, limiting practical use in real-world applications. Translating these task instructions into caption-style prompts for fixed text-to-image diffusion models is particularly challenging. On the other hand, specialized models exist for specific aspects (Fig. 1) of SubRep, like local inpainting [75, 80, 38, 62, 13], subject completion [76], and local harmonization [69, 77, 64]. However, combining components from these models can make the SubRep system bulky and less elegant. Given the shared generative nature of these sub-tasks, our study raises an intriguing question: "Can we achieve all these sub-tasks using a single model?"
@@
To answer this question, we introduce "task inversion", a novel concept that learns latent embeddings as alternative of text conditions to guide diffusion models with specific task instructions. The embedding space of text prompts in diffusion models offers versatility beyond just captions. Employing prompt tuning at the task level allows us to learn latent embeddings to guide diffusion models based on task instructions. Task inversion enables diffusion models to adapt to various tasks by adjusting task-level "text" prompts. Unlike textual inversion [17] which learns image-dependent caption prompts and prompt tuning [35, 42] which learns domain adaptation, our method employs task-level instructional prompts to approximate optimal text prompts for each image in a specific task, transforming text-to-image diffusion model into task-to-image model. Our approach pioneers the systematic use of learned embeddings across various generative sub-tasks within a single SD, effectively addressing the complex challenge of SubRep.
@@
To formally address the SubRep task, we propose the SEgment-gEnerate-and-bLEnd (SEELE) framework. As in Fig. 2, SEELE manages the subject repositioning with a pre-processing, manipulation, post-processing pipeline. i) In the pre-processing stage, SEELE segments the subject based on user-specified points, bounding boxes, or text prompts. With the provided moving direction, SEELE relocates the subject while considering occlusion relationships between subjects. ii) In the manipulation stage, SEELE uses a single diffusion model guided by learned task prompts to handle subject removal and completion. iii) In the post-processing stage, SEELE harmonizes the repositioned subject to blend seamlessly with adjacent regions.
@@
We’ve curated a dataset named ReS to test subject repositioning algorithms in real-world scenarios. We made efforts in covering various scenes and times to give a wide range of examples. Particularly, the real-world images for this task demand very exhaustively ground-truth annotation, including the mask of the repositioned subject and the moving direction. We annotate the mask using SAM [33] and manual refinement, and estimating the moving direction based on the center point of masks in the paired image. Additionally, we also provide amodal masks for subjects that are partly hidden. This results paired real image, actually diverse enough to support the evaluation of our task, as illustrated in Fig. 2(b). As far as we know, this is the first dataset designed specifically for subject repositioning. It’s diverse and well-organized, making it a great benchmark for validating methods for this task.
Contributions. Our contributions are as follows:
@@
We introduce SEgment-gEnerate-and-bLEnd (SEELE) framework, addressing multiple generative tasks with one diffusion model. Not only does it offer an application akin to Google’s magic editor, but it also organizes each subtask efficiently using a shared SD backbone. Furthermore, our approach provides additional features beyond the magic editor, including occlusion and perspective preservation, as well as local harmonization.
@@ 2 Subject Repositioning
Subject repositioning (SubRep) relocates the user-specified subject within an image. This seemingly simple task is actually challenging, requiring coordination of multiple sub-tasks and interaction between user and learning models.
User inputs. An illustration of the user inputs is shown in Fig. 2(a). SubRep follows user intention to identify the subject, move it to the desired location, complete it, and address disparities. Particularly, the user identifies the interested subject via pointing, bounding box, or text prompts. Then, the user provides the desired repositioning location via dragging or direction. The system further requires the user to indicate the occluded part of the subject for completion, and whether to run postprocessing algorithms for minimizing visual differences.
ReS dataset. To evaluate the effectiveness of subject repositioning algorithms, we curated a benchmark dataset called ReS. It includes paired images, each with dimensions 40323024: one image features a repositioned subject while the other elements remain constant. These images were collected from over 20 indoor and outdoor scenes, featuring subjects from over 50 categories. This diversity enables effective simulation of real-world applications, making our dataset suitable for evaluating our SEELE model.
We also contribute very detailed annotations to this dataset. Particularly, The masks for the repositioned subjects were initially generated using SAM and refined by multiple experts. Occluded masks were provided for subject completion. The direction of repositioning was estimated by measuring the distance between the center points of the masks in each image pair. For each paired image in the dataset, we can assess subject repositioning performance from one image to the other and in reverse, resulting in double testing examples. Fig. 2(b) illustrates the ReS dataset. We release the ReS dataset at https://github.com/Yikai-Wang/ReS to encourage research in subject repositioning.
@@ 3 SEELE Framework for Subject Repositioning
Task decomposition. To tackle this task, we introduce the SEgment-gEnerate-and-bLEnd (SEELE) framework, shown in Fig. 2. Specifically, SEELE breaks down the task into three stages: preprocessing, manipulation, and post-processing. Preprocessing handles non-generative tasks, while manipulation and postprocessing require generative capabilities. We use a unified diffusion model for all generative sub-tasks and pre-trained models for non-generative tasks in SEELE.
i) The preprocessing addresses how to precisely locate the specified subject with minimal user input, considering that the subject may be a single object, part of an object, or a group of objects identified by the user’s intention; reposition the identified subject to the desired location; and also identify occlusion relationships to maintain geometric consistency. Additionally, adjusting the subject’s size might be necessary to maintain the perspective relationship.
@@
iii) The postprocessing stage focuses on minimizing visual differences between the repositioned subject and its new surroundings. This involves fixing inconsistencies in both appearance and geometry, including blending unnatural boundaries, aligning illumination statistics, and, at times, creating realistic shadows.
Pre-processing. For point and bounding box inputs for identifying subject, we utilize SAM [33] for user interaction and employ SAM-HQ [29] to enhance the quality of segmenting subjects with intricate structures. To enable text inputs, we follow SeMani [67] to indirectly implement a text-guided SAM mode. Specifically, we first employ SAM to segment the entire image into distinct subjects. Then we identify the most similar one using the mask-adapted CLIP [40].
@@
SEELE handles the potential occlusion between the moved subject and other elements in the image. If there are other subjects present at the desired location, SEELE employs the monocular depth estimation algorithm MiDaS [54] to discern occlusion relationships between subjects. SEELE will then appropriately mask the occluded portions of the subject if the user wants to preserve these occlusion relationships. MiDaS is also used to estimate the perspective relationships among subjects and resize the subject accordingly to maintain geometric consistency. For subjects with ambiguous boundaries, SEELE incorporates the ViTMatte matting algorithm [71] for better compositing with surrounding areas. An illustrated comparison of incorporated modules can be found in Fig. 8.
@@
Manipulation. In this stage, SEELE deals with the primary tasks of manipulating subjects, including subject removal and subject completion, as illustrated in Fig. 2. Critically, such two steps can be effectively solved by a single generative model, as the masked region of both steps should be filled in to match the surrounding areas. However, these two sub-tasks require different information and types of masks. Particularly, for subject removal, a non-semantic inpainting is applied uniformly from the unmasked regions, using a typical object-shaped mask. This often falsely results in the creation of new, random subjects within the holes. On the other hand, subject completion involves semantic-rich inpainting and aims to incorporate the majority of the masked region as part of the subject. Critically, to adapt the same diffusion model to the different generation directions needed for the above sub-tasks, we propose the task inversion technique in SEELE. This technique guides the diffusion model according to specific task instructions. Thus, with the learned remove-prompt and complete-prompt, SEELE tackles these sub-tasks via a single generative model. An illustrated comparison between different task-prompts can be found in Fig. 6(a).
@@
Post-processing. In the final stage, SEELE harmoniously blends the repositioned subject with its surroundings by tackling two challenges below. The illustrated comparison of post-processing can be found in Fig. 8.
@@
i) Local harmonization ensures natural appearance in boundary and lighting statistics. SEELE confines this process to the relocated subject to avoid affecting other image parts. It takes the image and a mask indicating the subject’s repositioning as inputs. However, the stable diffusion model is initially trained to generate new concepts within the masked region, conflicting with our goal of only ensuring consistency in the masked region and its surroundings. To address this, SEELE adapts the model by learning a harmonize-prompt with LoRA adapter [22] to guide masked regions. This can also be integrated into the same diffusion model used in the manipulation stage with our newly proposed design.
@@
ii) Shadow generation aims to create realistic shadows for repositioned subjects, enhancing the realism. Generating high-fidelity shadows in high-resolution images of diverse subjects remains challenging. SEELE uses the diffusion model for shadow generation, addressing two scenarios: 1) If the subject already has shadows, we use complete-prompt for shadow completion. 2) For subjects without shadows, we follow user-intention to locate the desired shadow area. This task then transforms into a local harmonization process for lighting.
@@ 3.1 Task Inversion
Subject removal fills the void in original area without creating new subjects;
Subject completion completes the repositioned subject within masked region;
Subject harmonization blends subject without inducing new elements.
@@
These requirements lead to different generation paths. However, our goal is to adapt frozen text-to-image diffusion inpainting models for all of these sub-tasks.
To address this challenge, we introduce task inversion, training prompts to guide the diffusion model while keeping the backbone fixed. Instead of traditional text prompts, we utilize the adaptable representations acting as instruction prompts, such as “complete the subject”. The challenge lies in the domain gap where text-to-image diffusion model is not trained from instruction prompts. Our experiments show that compared with unconditional generation and simple semantic and instructional text prompt-guided generation, the learned task prompts significantly improves the inpainting model in standard inpainting and outpainting tasks (see Tab. 2), as well as sub-tasks of subject repositioning (see Tab. 1). Thus our learned task prompts can be used as an alternative of image-dependent text prompt for subject repositioning to minimize user effort. Furthermore, task inversion allows the integration of different generative sub-tasks for subject repositioning using stable diffusion. This integration avoids the need for introducing new generative models or adding extensive modules or parameters, highlighting the plug-and-play nature of task inversion.
@@
Task inversion adheres to the original training objectives of diffusion models. Specifically, denote the training image as , the local mask as , the learnable task prompt as . Our objective is
@@
where is the random noise; is the diffusion model, is the normalized noise-level; is the noised image, is element-wise multiplication; and is the Frobenius norm. When training with Eq. (1), the and the conditioning model is frozen, making the embedding the only learnable parameters.
@@
Our task inversion is a distinctive approach, influenced by various existing works but with clear differences. The instruction prompt mentioned for our task inversion goes beyond the training data’s scope, where the text describes the content of image, potentially affecting the desired generation results in practice. Recent advancements in textual inversion [17] emphasize the potential to comprehend user-specified concepts within the embedding space. In contrast, prompt tuning [35, 42] enhances adaptation to specific domains by introducing learnable tokens to the inputs. Unlike textual inversion, which trains a few tokens for visual understanding, our task inversion trains the whole latent to provide task instruction. Our task inversion differs prompt-tuning in that: prompt-tuning adds new tokens, while our approach replaces text condition inputs. We don’t depend on text inputs to guide the diffusion model. See Fig. 3(a) for the distinction.

### v3

# Repositioning the Subject within Image

@@ Abstract
Current image manipulation primarily centers on static manipulation, such as replacing specific regions within an image or altering its overall style. In this paper, we introduce an innovative dynamic manipulation task, subject repositioning. This task involves relocating a user-specified subject to a desired position while preserving the image’s fidelity. Our research reveals that the fundamental sub-tasks of subject repositioning, which include filling the void left by the repositioned subject, reconstructing obscured portions of the subject and blending the subject to be consistent with surrounding areas, can be effectively reformulated as a unified, prompt-guided inpainting task. Consequently, we can employ a single diffusion generative model to address these sub-tasks using various task prompts learned through our proposed task inversion technique. Additionally, we integrate pre-processing and post-processing techniques to further enhance the quality of subject repositioning. These elements together form our SEgment-gEnerate-and-bLEnd (SEELE) framework. To assess SEELE’s effectiveness in subject repositioning, we assemble a real-world subject repositioning dataset called ReS. Results of SEELE on ReS demonstrate its efficacy. Code and ReS dataset are available at https://yikai-wang.github.io/seele/.
@@ 1 Introduction
In 2023, Google Photos introduced an AI editing feature allowing users to reposition subjects within their images (Google, 2023). However, a lack of technical documentation limits understanding of this feature. Some researches have touched on aspects of it. Iizuka et al. (2014) explored object repositioning before the deep learning era, using user inputs like ground regions and bounding boxes. In the deep learning era, fields like scene decomposition (Zheng et al., 2021) and de-occlusion (Zhan et al., 2020) enable manipulation of object positions after the explicit understanding of scene and object relationships. This paper addresses general Subject Repositioning (SubRep) task without explicit scene understanding. Our aim is to address SubRep via a meticulously crafted solution, driven by a single diffusion model.
@@
From an academic perspective, this task falls under image manipulation (Gatys et al., 2016; Isola et al., 2017; Zhu et al., 2017; Wang et al., 2018; El-Nouby et al., 2019; Fu et al., 2020; Zhang et al., 2021). Recent advancements in large-scale generative models have fueled interest in this field. These models, including generative adversarial models (Goodfellow et al., 2014), variational autoencoders (Kingma & Welling, 2014), auto-regressive models (Vaswani et al., 2017), and notably, diffusion models (Sohl-Dickstein et al., 2015), demonstrate impressive image manipulation capabilities with expanding model architectures and training datasets (Rombach et al., 2022; Kawar et al., 2022; Chang et al., 2023). However, current image manipulation methods primarily target "static" alterations, modifying specific image regions using cues like natural language, sketches, or layouts (El-Nouby et al., 2019; Zhang et al., 2021; Fu et al., 2020). Another aspect involves style-transfer tasks, transforming overall image styles such as converting photos into anime pictures or paintings (Chen et al., 2018; Wang et al., 2018; Jiang et al., 2021). Some extend to video manipulation, altering style or subjects over time (Kim et al., 2019; Xu et al., 2019; Fu et al., 2022). In contrast, subject repositioning dynamically relocates selected subjects within a single image while leaving the rest unchanged.
@@
The SubRep task involves multiple stages, including non-generative and generative tasks. Existing pre-trained models are effective for non-generative tasks like segmenting subjects (Kirillov et al., 2023) and estimating occlusion relationships (Ranftl et al., 2020). Our focus lies on the generative tasks of SubRep, including: i) Subject removal: The generative model must fill voids left after repositioning without introducing new elements. ii) Subject completion: If the repositioned subject is partially obscured, the model must complete it to maintain integrity. iii) Subject harmonization: The repositioned subject should blend with surrounding areas. All these sub-tasks demand unique generative capabilities.
@@
The most powerful text-to-image diffusion models (Nichol et al., 2022; Ho et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022) show potential promise for SubRep. However, a key challenge is finding suitable text prompts, as these models are usually trained with image captions rather than task-specific instructions. The best prompts are often image-dependent and hard to generalize, limiting practical use in real-world applications. Translating these task instructions into caption-style prompts for fixed text-to-image diffusion models is particularly challenging. On the other hand, specialized models exist for specific aspects (Figure 1) of SubRep, like local inpainting (Zeng et al., 2020; Zhao et al., 2021; Li et al., 2022; Suvorov et al., 2022; Dong et al., 2022), subject completion (Zhan et al., 2020), and local harmonization (Xu et al., 2017; Zhang et al., 2020; Tsai et al., 2017). However, combining components from these models can make the SubRep system bulky and less elegant. Given the shared generative nature of these sub-tasks, our study raises an intriguing question: "Can we achieve all these sub-tasks using a single model?"
@@
To answer this question, we introduce "task inversion", a novel concept that learns latent embeddings as alternative of text conditions to guide diffusion models with specific task instructions. The embedding space of text prompts in diffusion models offers versatility beyond just captions. Employing prompt tuning at the task level allows us to learn latent embeddings to guide diffusion models based on task instructions. Task inversion enables diffusion models to adapt to various tasks by adjusting task-level "text" prompts. Unlike textual inversion (Gal et al., 2022) which learns image-dependent caption prompts and prompt tuning (Lester et al., 2021; Liu et al., 2021a) which learns domain adaptation, our method employs task-level instructional prompts to approximate optimal text prompts for each image in a specific task, transforming text-to-image diffusion model into task-to-image model. Our approach pioneers the systematic use of learned embeddings across generative sub-tasks within a single SD, effectively addressing the complex challenge of SubRep.
@@
To formally address the SubRep task, we propose the SEgment-gEnerate-and-bLEnd (SEELE) framework. As in Figure 2, SEELE manages the subject repositioning with a pre-processing, manipulation, post-processing pipeline. i) In the pre-processing stage, SEELE segments the subject based on user-specified points, bounding boxes, or text prompts. With the provided moving direction, SEELE relocates the subject while considering occlusion relationships between subjects. ii) In the manipulation stage, SEELE uses a single diffusion model guided by learned task prompts to handle subject removal and completion. iii) In the post-processing stage, SEELE harmonizes the repositioned subject to blend with adjacent regions.
@@
We’ve curated a dataset named ReS to test subject repositioning algorithms in real-world scenarios. We made efforts in covering various scenes and times to give a wide range of examples. Particularly, the real-world images for this task demand very exhaustively ground-truth annotation, including the mask of the repositioned subject and the moving direction. We annotate the mask using SAM (Kirillov et al., 2023) and manual refinement, and estimating the moving direction based on the center point of masks in the paired image. Additionally, we also provide amodal masks for subjects that are partly hidden. This results paired real image, actually diverse enough to support the evaluation of our task, as illustrated in Figure LABEL:fig:relocateobj. As far as we know, this is the first dataset designed specifically for subject repositioning. It’s diverse and well-organized, making it a great benchmark for validating methods for this task.
Contributions Our contributions are as follows:
@@
We present the SEgment-gEnerate-and-bLEnd (SEELE) framework, which tackles various generative tasks using a single diffusion model. SEELE offers an application akin to Google’s magic editor. Additionally, SEELE goes beyond the Magic Editor by offering advanced features like preserving occlusion and perspective, as well as local harmonization.
@@ 2 Subject Repositioning
Subject repositioning (SubRep) relocates the user-specified subject within an image. The "subject" can be anything the user focuses on, such as a part of an object, an entire object, or multiple objects. Although it sounds straightforward, this task is quite complex. It requires coordination of multiple sub-tasks and interaction between user and learning models.
User inputs An illustration of the user inputs is shown in Figure LABEL:fig:user-effort. SubRep follows user intention to identify the subject, move it to the desired location, complete it, and address disparities. Particularly, the user identifies the interested subject via pointing, bounding box, or text prompts. Then, the user provides the desired repositioning location via dragging or direction. The user also needs to specify the occluded part of the subject for completion, and decide whether to apply post-processing to reduce visible differences.
ReS dataset To evaluate the effectiveness of subject repositioning algorithms, we curated a benchmark dataset called ReS. It includes paired images: one image features a repositioned subject while the other elements remain constant. These images were collected from over 20 indoor and outdoor scenes, featuring subjects from over 50 categories. This diversity enables effective simulation of real-world applications, making our dataset suitable for evaluating our SEELE model.
We also contribute very detailed annotations to this dataset. Particularly, The masks for the repositioned subjects were initially generated using SAM and refined by multiple experts. Occluded masks were provided for subject completion. The direction of repositioning was estimated by measuring the distance between the center points of the masks in each image pair. For each paired image in the dataset, we can assess subject repositioning performance from one image to the other and in reverse, resulting in double testing examples. Figure LABEL:fig:relocateobj illustrates the ReS dataset. We release the ReS dataset at https://yikai-wang.github.io/seele/ to encourage research in subject repositioning.
@@ 3 SEELE Framework for Subject Repositioning
Task decomposition To tackle this task, we introduce the SEgment-gEnerate-and-bLEnd (SEELE) framework, shown in Figure 2. Specifically, SEELE breaks down the task into three stages: pre-processing, manipulation, and post-processing. Pre-processing handles non-generative tasks, while manipulation and post-processing require generative capabilities. We use a unified diffusion model for all generative sub-tasks and pre-trained models for non-generative tasks in SEELE.
i) The pre-processing addresses how to precisely locate the specified subject with minimal user input, considering that the subject may be a single object, part of an object, or a group of objects identified by the user’s intention; reposition the identified subject to the desired location; and also identify occlusion relationships to maintain geometric consistency. Additionally, adjusting the subject’s size might be necessary to maintain the perspective relationship.
@@
iii) The post-processing stage focuses on minimizing visual differences between the repositioned subject and its new surroundings. This involves fixing inconsistencies in both appearance and geometry, including blending unnatural boundaries, aligning illumination statistics, and, at times, creating realistic shadows.
Pre-processing For point and bounding box inputs for identifying subject, we utilize SAM (Kirillov et al., 2023) for user interaction and employ SAM-HQ (Ke et al., 2023) to enhance the quality of segmenting subjects with intricate structures. To enable text inputs, we follow SeMani (Wang et al., 2023) to indirectly implement a text-guided SAM mode. Specifically, we first employ SAM to segment the entire image into distinct subjects. Then we identify the most similar one using the mask-adapted CLIP (Liang et al., 2022).
@@
SEELE handles the potential occlusion between the moved subject and other elements in the image. If there are other subjects present at the desired location, SEELE employs the monocular depth estimation algorithm MiDaS (Ranftl et al., 2020) to discern occlusion relationships between subjects. SEELE will then appropriately mask the occluded portions of the subject if the user wants to preserve these occlusion relationships. MiDaS is also used to estimate the perspective relationships among subjects and resize the subject accordingly to maintain geometric consistency. For subjects with ambiguous boundaries, SEELE incorporates the ViTMatte matting algorithm (Yao et al., 2023) for better compositing with surrounding areas. An illustrated comparison of incorporated modules can be found in Figure 7.
@@
Manipulation In this stage, SEELE deals with the primary tasks of manipulating subjects, including subject removal and subject completion, as illustrated in Figure 2. Critically, such two steps can be effectively solved by a single generative model, as the masked region of both steps should be filled in to match the surrounding areas. However, these two sub-tasks require different information and types of masks. Particularly, for subject removal, a non-semantic inpainting is applied uniformly from the unmasked regions, using a typical object-shaped mask. This often falsely results in the creation of new, random subjects within the holes. On the other hand, subject completion involves semantic-rich inpainting and aims to incorporate the majority of the masked region as part of the subject. Critically, to adapt the same diffusion model to the different generation directions needed for the above sub-tasks, we propose the task inversion technique in SEELE. This technique guides the diffusion model according to specific task instructions. Thus, with the learned remove-prompt and complete-prompt, SEELE tackles these sub-tasks via a single generative model. An illustrated comparison between different task-prompts can be found in Figure LABEL:fig:inpaint-vs-complete.
@@
Post-processing In the final stage, SEELE blends the repositioned subject with its surroundings by tackling two challenges below. The illustrated comparison of post-processing can be found in Figure 7.
@@
i) Local harmonization ensures natural appearance in boundary and lighting statistics. SEELE confines this process to the relocated subject to avoid affecting other image parts. It takes the image and a mask indicating the subject’s repositioning as inputs. However, the stable diffusion model is initially trained to generate new contents within the masked region, conflicting with our goal of only ensuring consistency in the masked region and its surroundings. To address this, SEELE adapts the model by learning a harmonize-prompt with LoRA adapter (Hu et al., 2021) to guide masked regions. This can also be integrated into the same diffusion model used in the manipulation stage with our newly proposed design.
@@
ii) Shadow generation aims to create realistic shadows for repositioned subjects, enhancing the realism. Generating high-fidelity shadows in high-resolution images of diverse subjects remains challenging. SEELE uses the diffusion model for shadow generation, addressing two scenarios: 1) If the subject already has shadows, we use complete-prompt for shadow completion. 2) For subjects without shadows, we follow user-intention to locate the desired shadow area. This task then transforms into a local harmonization process.
@@ 3.1 Task Inversion
1) Subject removal fills the void in original area without creating new subjects;
2) Subject completion completes the repositioned subject within masked region;
3) Subject harmonization blends subject without inducing new elements.
@@
These requirements lead to different generation paths. Our goal is to adapt frozen text-to-image diffusion inpainting models for all of these sub-tasks.
Task prompts To address these challenges, we introduce task inversion, a method that trains prompts to guide the diffusion model for specific generation tasks while keeping its backbone fixed.
In standard text-to-image diffusion models, text prompts like "a cute cat" are processed through a text encoder, which generates token sequences to guide image generation. Task inversion eliminates the need for text inputs and the text encoder. Instead, we train learnable prompts, called task prompts, to serve as input sequences. These task prompts directly guide the model to perform specific tasks. Conceptually, they act as instructions such as "complete the subject". See Figure LABEL:fig:ti-vs-ti for an illustration.
A key challenge is the domain gap: text-to-image diffusion models are not originally trained to respond to instruction-based prompts. However, our experiments demonstrate that learned task prompts significantly enhance performance. Compared to unconditional generation or simple semantic and instructional text prompts, task prompts deliver substantial improvements in standard inpainting and outpainting tasks (see Table 2) and sub-tasks like subject repositioning (see Table 1).
Our approach also reduces user effort by serving as an alternative to image-dependent text prompts for subject repositioning. Moreover, task inversion seamlessly integrates various generative sub-tasks for subject repositioning using stable diffusion. This eliminates the need for new generative models or extensive additional modules, emphasizing its plug-and-play simplicity.
Inverse to learn task prompt We aim to learn an optimal task prompt that conditions diffusion models for specific inpainting tasks. This task prompt is trained on input-output pairs, enabling it to translate task instructions from the training dataset into learned representations.
In text-to-image diffusion inpainting models, input conditions are typically text strings embedded into a sequence of vectors. For instance, in SD 2.0, a text encoder processes the text into a sequence of size , which is then passed through the U-Net’s cross-attention layers. Instead of using a text-based approach, our method directly learns a sequence of size to represent the task prompts.
Unlike user-driven prompts for specifying object location or direction, these task prompts are pre-learned during training. Each prompt is associated with specific input-output pairs tailored for the task, as explained in Sec. 3.2. During inference, task prompts are integrated into the pipeline shown in Figure 2. Users can then select between subject completion or harmonization via a simple button.
@@
Formally, task inversion adheres to the original training objectives of diffusion models. Specifically, denote the training image as , the local mask as , the learnable task prompt as . Our objective is
@@
where is the random noise; is the diffusion model, is the normalized noise-level; is the noised image, is element-wise multiplication; and is the Frobenius norm. When training with Eq. (1), the is frozen, making the embedding the only learnable parameters.
@@
Our task inversion is a distinctive approach, influenced by various existing works but with clear differences. The instruction prompt mentioned for our task inversion goes beyond the training data’s scope, where the text describes the content of image, potentially affecting the desired generation results in practice. Recent advancements in textual inversion (Gal et al., 2022) emphasize the potential to comprehend user-specified concepts within the embedding space. In contrast, prompt tuning (Lester et al., 2021; Liu et al., 2021a) enhances adaptation to specific domains by introducing learnable tokens to the inputs. Unlike textual inversion, which trains a few tokens for visual understanding, our task inversion trains the whole latent to provide task instruction. Our task inversion differs prompt-tuning in that: prompt-tuning adds new tokens, while our approach replaces text condition inputs. We don’t depend on text inputs to guide the diffusion model. See Figure LABEL:fig:ti-vs-ti for the distinction.